OpenStudy (anonymous):

integral of root (1+sin2x)dx ..??? hint pll

OpenStudy (luigi0210):

\[\int\limits \sqrt{1+\sin^2x} dx\]

OpenStudy (anonymous):

nooo

OpenStudy (luigi0210):

Is it just sin 2x?

OpenStudy (anonymous):

\[\int\limits \sqrt{1+\sin2x} dx\]

OpenStudy (luigi0210):

.. Well we could start by rewriting it.. \[\int\limits (1+\sin2x)^{1/2} dx\]

OpenStudy (anonymous):

divide and multiply by sqrt ( 1 - sin2x )

OpenStudy (anonymous):